OpenStudy (anonymous):

Find a quadratic model for the set of values: (-2, -20), (0, -4), (4, -20). Show your work

OpenStudy (sdfgsdfgs):

Since u need to show ur work....the general quadratic model is y=ax^2 + bx + c Now we need to find the constants a, b, c ok?

OpenStudy (anonymous):

How do we do that

OpenStudy (sdfgsdfgs):

U are given 3 points: (-2, -20), (0, -4), (4, -20) So u can use them to solve for abc

OpenStudy (anonymous):

I suck at this and I don't get it

OpenStudy (sdfgsdfgs):

I will do the first one for u, @Raylynngrace :) y=ax^2 + bx + c for (-2, -20) so -20 = a*(-2)^2 + b*(-2) + c -20 = 4a - 2b + c ok?

OpenStudy (anonymous):

Okay I get that now lol

OpenStudy (sdfgsdfgs):

Cool! Show me how u set up with the other two given points plz?

OpenStudy (anonymous):

idk if I did this right lol but for the next one I got y=ax^2+bx+c -4=a*(0)^2+b*(0) -4=a-b+c

OpenStudy (sdfgsdfgs):

good try but u missed something...it should be like -4=a*(0)^2+b*(0) + c -4=0+0+c c=-4 hey we found c already!

OpenStudy (anonymous):

Least I tried lol can you help me with the last one

OpenStudy (sdfgsdfgs):

nope u do it plz? i will help of cus :)

OpenStudy (anonymous):

y=ax^2+bx+c -20=a*(4)^2+b*(4) -20=16a+4b

OpenStudy (anonymous):

Did I get this one wrong to @sdfgsdfgs

OpenStudy (sdfgsdfgs):

@raylynngrace hahahaa u keep forgetting the c term....so it should be: -20=a*(4)^2+b*(4)+c -20=16a+4b+c now u got c=-4 already and from the first point: -20 = 4a - 2b + c U can solve for a and b :)

OpenStudy (anonymous):

@Raylynngrace This will help you a bit. Try !! http://www.softschools.com/math/algebra/quadratic_functions/quadratic_function_with_three_points/

OpenStudy (anonymous):

Thank you both for your help @sdfgsdfgs @JoshKoikkara